2x+y=5
2x+y=1 substatution

Answer by mananth(16946)

2 x + 1 y = 5 .............1
1 y = 1 -2 x ---------2
/ 1 =
y = 1 -2 x
Substitute y in (1)
2 x + 1 ( 1 -2 x ) =
2 x + 1 + -2 x = 5
2 x -2 x = 5 -1
0 x = 4
/ 0
x= undefined
So the system has no solution. The lines are parallel